# RAS: Retrieval-And-Structuring for Knowledge-Intensive LLM Generation

### v2

@@ Abstract
Large language models (LLMs) have achieved impressive performance on knowledge-intensive tasks, yet they often struggle with multi-step reasoning due to the unstructured nature of retrieved context. While retrieval-augmented generation (RAG) methods provide external information, the lack of explicit organization among retrieved passages limits their effectiveness, leading to brittle reasoning pathways. Recent interpretability studies highlighting the importance of structured intermediate reasoning further align with this perspective. We propose Retrieval-And-Structuring (RAS), a framework that dynamically constructs query-specific knowledge graphs through iterative retrieval and structured knowledge building. RAS interleaves targeted retrieval planning with incremental graph construction, enabling models to assemble and reason over evolving knowledge structures tailored to each query. On seven knowledge-intensive benchmarks, RAS consistently outperforms strong baselines, achieving up to 6.4% and 7.0% gains with open-source and proprietary LLMs, respectively. Our results demonstrate that dynamic, query-specific knowledge structuring offers a robust path to improving reasoning accuracy and robustness in language model generation. Our data and code can be found at https://github.com/pat-jj/RAS.
@@ 1 Introduction
Complex reasoning tasks such as scientific analysis or multi-hop question answering demand both comprehensive knowledge and structured logical thinking [1]. While large language models (LLMs) have achieved remarkable performance across a wide range of natural language processing tasks [2, 3, 4], they often struggle with knowledge-intensive reasoning due to the absence of precise, logically organized information [5, 6]. This limitation has motivated growing research into augmenting LLMs with structured knowledge to enhance their reasoning capabilities [7].
Retrieval-augmented generation (RAG) approaches have emerged as a promising direction by providing LLMs with additional context from retrieved passages [8, 9, 10, 11]. However, these methods frequently face hallucination challenges [12, 13], where generated content deviates from retrieved information. This issue stems primarily from the unstructured nature of retrieved passages, requiring the model to implicitly bridge gaps during generation. Recent research in model interpretability further highlights this issue [14]. Attribution graph analyses have revealed that LLMs internally attempt multi-hop reasoning by chaining facts across different parts of context, but failures in these reasoning pathways are correlated with hallucinations and factual inconsistencies. These insights suggest that explicitly providing structured intermediate knowledge could significantly enhance the reliability and transparency of LLM reasoning. Recent efforts have explored integrating knowledge graphs (KGs) with LLMs [15, 16, 11, 17] to provide compact, relational representations that enable more interpretable reasoning [18, 19, 20]. However, existing approaches largely rely on static, corpus-wide graphs, which may not adapt well to the specific reasoning needs of individual queries. Constructing and maintaining large-scale KGs also incurs substantial costs in open-domain settings.
@@
To address these challenges, we propose Retrieval-And-Structuring (RAS), a framework that dynamically constructs and reasons over query-specific knowledge graphs through iterative retrieval and structured knowledge building. RAS interleaves three key components: (1) a planning module that strategically identifies knowledge gaps and generates targeted sub-queries, (2) a retrieval-and-structuring module that extracts factual triples from retrieved passages and incrementally builds a query-specific graph, and (3) a knowledge-augmented answering module that generates outputs conditioned on the accumulated structured knowledge.
@@
RAS addresses several critical limitations of prior methods. Unlike traditional RAG approaches that rely on single-pass retrieval [8, 9, 10], RAS employs iterative planning to actively identify and fill knowledge gaps during inference. In contrast to static KG-based approaches [11, 17], RAS dynamically constructs query-specific knowledge graphs tailored to the evolving reasoning needs of each query. This targeted construction captures only task-relevant information, avoiding the inefficiencies and noise associated with static, corpus-level graphs, which enables more precise and robust language generation.
@@
Through extensive evaluations across seven benchmarks spanning open-domain QA, closed-set QA, and long-form generation, RAS consistently outperforms strong baselines by 6.4% with open-source LLMs and 7.0% with proprietary models. Our main contributions are:
We introduce RAS, a novel framework for dynamic knowledge structuring through iterative retrieval and graph construction, tailored to each query.
@@
We develop a structure-aware planning and answering model that operates over evolving knowledge graphs.
We demonstrate RAS’s effectiveness through comprehensive experiments and ablation studies across diverse knowledge-intensive tasks.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG) [8, 9] enhances language model performance on knowledge-intensive tasks by incorporating retrieved passages into the model input, improving factual accuracy and grounding. Early RAG approaches retrieved a fixed number of passages once before generation [21, 22, 23], while later methods explored adaptive retrieval [24], where models dynamically issue retrieval queries based on uncertainty signals, or retrieval evaluation [25] to improve relevance. However, these techniques often struggle with complex or multi-hop reasoning due to their linear designs. Recent advances aim to strengthen retrieval quality by iterative refinement. Iterative retrieval-generation approaches [21, 26] leverage historical context to guide subsequent retrievals, while targeted retrieval strategies based on subqueries [27] or follow-up questions [28, 29] progressively enrich the retrieved evidence. Self-RAG [30] introduced self-reflective retrieval, allowing LLMs to selectively query for missing information and evaluate the quality of retrievals. RPG [31] further refined retrieval by extracting fine-grained paragraphs to enhance relevance. Despite these advances, existing methods face persistent challenges: retrieved context often contains redundant or irrelevant information, or conversely, discards essential facts during refinement. To address these issues, we propose converting retrieved content into a structured graph format and maintaining an iteratively enriched, query-specific knowledge graph, allowing the model to reason over an organized and progressively expanded knowledge structure.
Graph as Context for LLMs. Integrating structured graphs into language model workflows has emerged as a powerful approach for enhancing reasoning and factual accuracy [32, 33, 34, 35, 36, 37]. Graphs provide explicit, relational representations that help models go beyond flat text by making entity connections and multi-hop relationships more tractable. Recent work, such as GraphToken [38], shows that LLMs can process serialized graph inputs through token-level prompts [39] without modifying model parameters. G-Retriever [11] leverages global knowledge graphs (KGs) to retrieve entity-centric subgraphs, while GraphRAG-style methods [17, 40] construct corpus-level graphs and generate summaries to enhance retrieval quality. Despite their benefits, these methods rely heavily on static, global graphs, which often include irrelevant or redundant information and cannot adapt to the needs of individual queries. Moreover, building and maintaining large-scale KGs is computationally expensive, particularly in open-domain settings. This motivates a shift toward dynamic, inference-time graph construction. Our approach embraces this direction by proposing a retrieval-and-structuring framework that builds query-specific knowledge graphs on the fly. Rather than relying on fixed knowledge coverage, we iteratively retrieve and structure information into task-relevant graphs that evolve with the reasoning process. This dynamic graph construction aligns with recent interpretability findings [14], which suggest that LLM failures often stem from breakdowns in implicit reasoning chains—failures that explicit graph structuring can help mitigate.
@@ 3 Retrieval-And-Structuring (RAS) Framework
Effective knowledge-intensive language generation requires not only retrieving relevant information, but also structuring and reasoning over it systematically. We introduce Retrieval-And-Structuring (RAS), a framework that interleaves iterative retrieval planning with dynamic knowledge graph construction, enabling large language models (LLMs) to reason over progressively organized knowledge tailored to each query. Figure 1 illustrates the overall workflow.
Key Definitions. We define the core concepts used in the RAS framework as follows. The Main Query () denotes the original task input. A Subquery () is a focused retrieval query generated at iteration to obtain supporting evidence. Retrieved Passages () are the top- documents selected for each subquery. A text-to-triples model converts retrieved text into Extracted Triples (), structured as subject-predicate-object facts. These triples are incrementally accumulated into an evolving Knowledge Graph (), representing organized evidence about . The model produces an Action Plan () at each step, determining whether to continue retrieval ([SUBQ]), stop retrieval ([SUFFICIENT]), or, initially, answer directly without retrieval ([NO_RETRIEVAL]).
3.1 Knowledge State-Aware Planning
The Planning stage initiates and controls the retrieval-and-structuring process by dynamically assessing the current knowledge state.
@@ 3.1 Knowledge State-Aware Planning
[SUBQ] : If assesses that the query cannot be satisfactorily answered with its own knowledge, we start the iteration with the main query as the initial subquery, and move to the next stage (§3.2).
@@
[SUFFICIENT]: The accumulated knowledge is deemed sufficient to comprehensively address the main query . The iterative retrieval process terminates, and the framework proceeds to the Answering stage (§3.3).
The Planning stage serves as a key driver of the RAS framework’s iterative retrieval and refinement process. By dynamically assessing the adequacy of the retrieved knowledge and generating targeted sub-queries, it enables the efficient acquisition of query-relevant information.
3.2 Text Retrieval and Knowledge Structuring
This stage transforms retrieved text into structured knowledge and progressively builds the query-specific graph .
@@ 3.2 Text Retrieval and Knowledge Structuring
We use a standard dense retriever by default but note that RAS is compatible with more advanced information retrieval methods [41, 42, 43].
Text-to-Triples (Graph) Conversion. To extract essential factual information from the retrieved passages , we employ a text-to-triples model . This model is trained on the full WikiOfGraph dataset [44], which is a high-quality, LLM-curated text-to-triples corpus. Details of the training process are provided in Appendix C.1. The model generates structured triples in the following format:
where each triple represents a subject-predicate-object fact extracted from the text. This structured representation enables efficient downstream reasoning and facilitates integration with external knowledge graphs. Although is a lightweight LLM capable of fast inference using techniques such as quantization [45] and optimized inference frameworks like vLLM [46], the text-to-triples conversion can also be precomputed offline when maximal efficiency is required.
Iterative Knowledge Enrichment to Query-Specific KG. The extracted triples are then converted into a graph structure , where and denote the sets of nodes and edges, respectively. Each node corresponds to a unique subject or object entity in , while each edge represents a predicate connecting two entities. To enrich the graph with semantic information, the attributes of nodes and edges are obtained through a text encoder (Sentence-BERT [47]):
These semantic embeddings enable the model to capture the nuanced relationships between entities and facilitate reasoning over the knowledge graph.
To progressively enrich the query-related knowledge in response to the evolving sub-queries, the structured graph at each iteration is merged into an evolving KG specific to the main query :
After enriching with the new knowledge, the Planning module (§3.1) reassesses the current knowledge to determine the next step. Based on and the chain of previous subqueries and their associated graph information, the model decides whether to generate another focused subquery for additional retrieval or to proceed with answering (§3.3) if the accumulated knowledge is sufficient.
@@ 3.3 Knowledge-Augmented Answering
At the final stage, the model generates an answer to the main query either conditioned on knowledge graph and subquery chain when retrieval-and-structuring was processed, or directly when no retrieval is needed.
@@
Otherwise, after iterative knowledge enrichment concludes with [SUFFICIENT] plan, the answer is generated using encoded KG and subquery chain:
@@ 3.4 Sturcture-Aware Multitask Learning
The RAS framework is trained through a multitask setup that unifies knowledge planning and knowledge-augmented answering under a standard next-token prediction objective.
@@
Planning Task: Input: (encoded ), planning instruction , subquery-triples history, main query . Output: next plan .
Answering Task: Input: (encoded ), answering instruction , full subquery-triples history, main query . Output: final answer .
The model used in RAS is Graph LLM, an architecture adapted from prior work [38, 11]. It is trained to perform both planning and answering tasks via standard next-token prediction, conditioned on the instruction type and input structure. To the best of our knowledge, this is the first application of Graph LLMs to general language generation tasks such as open question answering.
@@ 4 Experiments
Training Data & Setting. We develop HotpotQA-SUBQ, a dataset constructed based on HotpotQA [1], to train our model’s action planning and answering capabilities. Our dataset creation begins with document filtering: using Claude-3.5-Sonnet [48], we identify and retain only the supporting documents necessary for answering the main query, removing irrelevant content. For each supporting document , we then iteratively generate a subquery , considering the main query, previous subqueries, and supporting documents. During iteration , when more supporting documents remain, we create training samples with input and output label “[SUBQ] ”, where represents triples extracted from document , and is the main query. For the final supporting document, we label the sample as “[SUFFICIENT]”. To identify queries that can be answered directly, we test our base LLM (LLaMA-2-7B) on HotpotQA’s main queries without context. For correctly answered queries, we create training samples with the main query as input and “[NO_RETRIEVAL]” as the output label. Additionally, to ensure fair comparison with existing approaches, we incorporate the subset of Arc-Easy (2,147 samples) and ASQA (3,897 samples) from Self-RAG’s training data, resulting in 208k training samples in total. We place detailed training data processing, dataset statistics, and data samples in Appendix B. For efficient inference, we train a text-to-triples model on the WikiOFGraph dataset [44] using LLaMA-3.2-3B-Instruct as the base model (see Appendix C), and deploy it using vLLM for optimized runtime performance. We present our hyperparameter study of each component in Appendix H.
Knowledge Sources. Due to the large sizes of wiki corpora, we employ faiss [49] for efficient vector searching over the dense index. Following the Self-RAG [30], we utilize the Wikipedia 2018 [50] by default, while specifically using the Wikipedia 2020 for PopQA to access more recent information. To optimize retrieval efficiency, we partition the index into five segments.

### v3

@@ Abstract
Large language models (LLMs) have achieved impressive performance on knowledge-intensive tasks, yet they often struggle with multi-step reasoning due to the unstructured nature of retrieved context. While retrieval-augmented generation (RAG) methods provide external information, the lack of explicit organization among retrieved passages limits their effectiveness, leading to brittle reasoning pathways. Recent interpretability studies highlighting the importance of structured intermediate reasoning further align with this perspective. We propose Retrieval-And-Structuring (RAS), a framework that dynamically constructs question-specific knowledge graphs through iterative retrieval and structured knowledge building. RAS interleaves targeted retrieval planning with incremental graph construction, enabling models to assemble and reason over evolving knowledge structures tailored to each query. On seven knowledge-intensive benchmarks, RAS consistently outperforms strong baselines, achieving up to 8.7% and 7.0% gains with proprietary and open-source LLMs, respectively. Our results demonstrate that dynamic, question-specific knowledge structuring offers a robust path to improving reasoning accuracy and robustness in language model generation.
@@ 1 Introduction
Complex reasoning tasks such as scientific analysis or multi-hop question answering demand both comprehensive knowledge and structured logical thinking (Yang et al., 2018). While large language models (LLMs) have achieved remarkable performance across a wide range of natural language processing tasks (Devlin et al., 2018; Brown et al., 2020), they often struggle with knowledge-intensive reasoning due to the absence of precise, logically organized information (Rae et al., 2021; Ling et al., 2024). This limitation has motivated growing research into augmenting LLMs with structured knowledge to enhance their reasoning capabilities (Wang et al., 2021).
Retrieval-augmented generation (RAG) approaches provide LLMs with additional context from retrieved passages (Guu et al., 2020; Lewis et al., 2020; Izacard and Grave, 2021; He et al., 2024), but often face hallucination challenges (Maynez et al., 2020; Zhang et al., 2023b), where generated content deviates from retrieved information. This issue stems from the unstructured nature of passages, which forces the model to implicitly bridge logical gaps. Briefly, interpretability analyses have suggested that LLMs attempt to chain facts across context internally, and failures in these implicit reasoning chains correlate with hallucinations (Lindsey et al., 2025). These findings reinforce the need for explicitly structured intermediate knowledge to guide reasoning.
Recent efforts have integrated knowledge graphs (KGs) with LLMs (Sun et al., 2019; Yu et al., 2022; He et al., 2024; Edge et al., 2024), providing compact relational representations that support more interpretable reasoning (Hogan et al., 2021; Jiang et al., 2024; Sun et al., 2023). However, existing approaches typically rely on static, corpus-wide graphs. This design introduces two limitations. First, global KGs are costly to build and maintain: indexing a corpus like Wikipedia 2018 can require millions of LLM calls and cost tens of thousands to millions of USD (see Appendix G). Second, global graphs often blend evidence from many documents, leading to ambiguous or even contradictory relations. For example, a global KG might simultaneously encode that Geoffrey Hinton is linked to deep learning, to cognitive neuroscience, and to critiques of large models—without clarifying which aspect is relevant to user query’s focus. Similarly, biomedical KGs may contain both positive and negative associations between a drug and a disease, reflecting conflicting studies. In contrast, a question-specific KG built from targeted documents resolves these conflicts by grounding relations in a coherent, query-relevant context.
@@
These limitations highlight the need for knowledge graphs that are constructed on demand, tailored to the query, and structured to support reasoning. To this end, we propose Retrieval-And-Structuring (RAS), a framework that dynamically constructs and reasons over question-specific knowledge graphs through iterative retrieval and structured knowledge building. The RAS process unfolds in three steps: (1) a planning step that identifies knowledge gaps and generates targeted sub-queries, (2) a retrieval-and-structuring step that extracts factual triples from retrieved passages and incrementally builds a question-specific graph, and (3) a knowledge-augmented answering step that produces final outputs conditioned on the accumulated structured knowledge.
@@
RAS addresses several limitations of prior methods. Unlike traditional RAG, which performs single-pass retrieval (Guu et al., 2020; Lewis et al., 2020; Izacard and Grave, 2021), RAS iteratively plans and fills knowledge gaps at inference. In contrast to static KG-based approaches (He et al., 2024; Edge et al., 2024), RAS dynamically constructs question-specific graphs tailored to each question, capturing only task-relevant information. This design avoids both the inefficiency of offline indexing and the noise of global graphs, enabling precise and robust reasoning.
@@
Through extensive evaluations across seven benchmarks spanning open-domain QA, closed-set QA, and long-form generation, RAS consistently outperforms strong baselines by 7.0% with open-source LLMs and 8.7% with proprietary models. Our main contributions are:
We propose RAS, a framework that dynamically builds question-specific knowledge graphs through iterative retrieval and structuring.
@@
We design a unified graph structure-aware model that jointly plans retrieval and generates answers over evolving knowledge graphs.
We show consistent gains across seven benchmarks, with up to 8.7% improvement over strong RAG baselines, while maintaining efficiency and scalability.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG). RAG enhances language model performance on knowledge-intensive tasks by incorporating retrieved passages into the model input (Guu et al., 2020; Lewis et al., 2020), improving factual accuracy and grounding. Early approaches retrieved a fixed number of passages once before generation (Shao et al., 2023; Es et al., 2024; Lyu et al., 2024a), while later methods explored adaptive retrieval (Jiang et al., 2023) or retrieval evaluation (Kim et al., 2024b) to improve relevance. Iterative retrieval-generation approaches (Shao et al., 2023; Guan et al., 2024) and targeted subquery strategies (Khattab et al., 2023; Yao et al., 2023; Press et al., 2022; Trivedi et al., 2023) progressively enrich the evidence. Self-RAG (Asai et al., 2023) introduced self-reflective retrieval, and RPG (Lyu et al., 2024b) extracted fine-grained paragraphs. More recent work (Jiang et al., 2025c; Mei et al., 2025) further improves RAG by reinforcement learning over search behaviors. Despite these advances, retrieved context often contains redundancy or misses critical facts. Our work departs from these by converting retrieved content into a structured, evolving graph aligned with the query.
Graph as Context for LLMs. Graphs offer explicit, relational structures that help models go beyond flat text by making multi-hop relationships more tractable (Yasunaga et al., 2021; 2022; Yu, 2022; Ju et al., 2022; Zhu et al., 2024; Gutiérrez et al., 2024). GraphToken (Perozzi et al., 2024) shows LLMs can process serialized graphs directly (Liu et al., 2021). G-Retriever (He et al., 2024) leverages global KGs for entity-centric subgraph retrieval, while GraphRAG-style methods (Edge et al., 2024; Jiang et al., 2025b) construct large corpus-level graphs with community summarization. These methods rely on static graphs, which are costly to construct (Appendix G) and often introduce irrelevant noise. By contrast, RAS builds question-specific knowledge graphs dynamically, eliminating prohibitive offline costs and providing denser, task-relevant context tailored to each reasoning trajectory. This design aligns with findings that many LLM errors stem from failed implicit reasoning chains (Lindsey et al., 2025), which explicit, query-focused structuring can mitigate.
@@ 3 Retrieval-And-Structuring (RAS) Framework
Effective knowledge-intensive language generation requires not only retrieving relevant information, but also structuring and reasoning over it systematically. We introduce Retrieval-And-Structuring (RAS), a framework that interleaves iterative retrieval planning with dynamic question-specific knowledge graph construction, enabling large language models (LLMs) to reason over progressively organized knowledge tailored to each query. Figure 1 illustrates the overall workflow.
Key Definitions. We define the core concepts used in the RAS framework as follows. The Main question () denotes the original task input. A subquery () is a focused retrieval query generated at iteration to obtain supporting evidence. At the first iteration, . Retrieved text () is a set of the top- documents retrieved by . A text-to-triples model converts retrieved text into triples (), structured as subject-predicate-object facts. These triples are incrementally accumulated into an evolving question-specific knowledge graph (), representing organized evidence related to . The model produces an plan () at each step, determining whether to continue retrieval ([SUBQ]), stop retrieval ([SUFFICIENT]), or, initially, answer directly without retrieval ([NO_RETRIEVAL]).
3.1 Knowledge-Aware Planning
The planning step initiates and controls the retrieval-and-structuring process by dynamically assessing the current knowledge state.
@@ 3.1 Knowledge-Aware Planning
[SUBQ] : If assesses that the query cannot be satisfactorily answered with its own knowledge, we start the iteration with the main question as the initial subquery, and move to the next stage (§3.2).
@@
[SUFFICIENT]: The accumulated knowledge is deemed sufficient to comprehensively address the main question . The iterative retrieval process terminates, and the framework proceeds to the Answering stage (§3.3).
Planning serves as a key driver of the RAS framework’s iterative retrieval and refinement process. By dynamically assessing the adequacy of the retrieved knowledge and generating targeted sub-queries, it enables the efficient acquisition of query-relevant information.
3.2 Text Retrieval and Structuring
Once [SUBQ] is detected, we use the subquery to retrieve the text and transform it into structured knowledge , which is progressively merged to the question-specific graph .
@@ 3.2 Text Retrieval and Structuring
We use a standard dense retriever by default but note that RAS is compatible with more advanced information retrieval methods (Chaudhary et al., 2023; Kang et al., 2024; Jiang et al., 2025a).
Text-to-Triples Conversion. To extract essential factual information from the retrieved passages , we employ a text-to-triples model . This model is trained on the full WikiOfGraph dataset (Kim et al., 2024a), which is a high-quality, LLM-curated text-to-triples corpus. Details of the training process are provided in Appendix D.1. The model generates structured triples in the following format:
where each triple represents a subject-predicate-object fact extracted from the text. This structured representation enables efficient downstream reasoning and facilitates integration with external knowledge graphs. Although is a lightweight LLM capable of fast inference using techniques such as quantization (Dettmers et al., 2022) and optimized inference frameworks like vLLM (Kwon et al., 2023), the text-to-triples conversion can be precomputed offline as well when maximal efficiency is required.
Iterative Knowledge Enrichment to Question-Specific KG. The extracted triples are then converted into a graph structure , where and denote the sets of nodes and edges, respectively. Each node corresponds to a unique subject or object entity in , while each edge represents a predicate connecting two entities. To enrich the graph with semantic information, the attributes of nodes and edges are obtained through Sentence-BERT (Reimers, 2019):
These semantic embeddings enable the model to capture the nuanced relationships between entities and facilitate reasoning over the KG.
To progressively enrich the question-related knowledge in response to the evolving sub-queries, the structured graph at each iteration is merged into an evolving KG specific to the main question :
After enriching with the new knowledge, we plan (§3.1) for the next step. Based on and the chain of previous subqueries and their associated graph information, the model decides whether to generate another focused subquery for additional retrieval or to proceed with answering (§3.3) if the accumulated knowledge is sufficient.
@@ 3.3 Knowledge-Augmented Answering
When answering is triggered, the model generates an answer to the main question either conditioned on knowledge graph and subquery chain when retrieval-and-structuring was processed, or directly when no retrieval is needed.
@@
Otherwise, after iterative knowledge enrichment concludes with [SUFFICIENT] plan or the maximum iteration is reached, the answer is generated using encoded KG and subquery chain:
@@ 3.4 Sturcture-Aware Multitask Learning
The RAS framework is trained through a multitask setup that unifies knowledge-aware planning and knowledge-augmented answering under a standard next-token prediction objective.
@@
Planning: Input: current encoded question-specific KG , planning instruction (shown in Figure 15), subquery-triples history (), main question . Output: next plan .
Answering: Input: final encoded question-specific KG , answering instruction (shown in Figure 16), subquery-triples history (), main question . Output: final answer .
@@ 4.1 Settings
Training Data & Setting. We develop HotpotQA-SUBQ, a dataset constructed based on HotpotQA (Yang et al., 2018), to train our model’s action planning and answering capabilities. Our dataset creation begins with document filtering: using Claude-3.5-Sonnet (Anthropic, 2024), we identify and retain only the supporting documents necessary for answering the main question, removing irrelevant content. For each supporting document , we then iteratively generate a subquery , considering the main question, previous subqueries, and supporting documents. During iteration , when more supporting documents remain, we create training samples with input and output label “[SUBQ] ”, where represents triples extracted from document , and is the main question. For the final supporting document, we label the sample as “[SUFFICIENT]”. To identify queries that can be answered directly, we test our base LLM (LLaMA-2-7B) on HotpotQA’s main queries without context. For correctly answered queries, we create training samples with the main question as input and “[NO_RETRIEVAL]” as the output label. Additionally, to ensure fair comparison with existing approaches, we incorporate the subset of Arc-Easy (2,147 samples) and ASQA (3,897 samples) from Self-RAG’s training data, resulting in 208k training samples in total. We place detailed training data processing, dataset statistics, and data samples in Appendix C. For efficient inference, we train a text-to-triples model on the WikiOFGraph dataset (Kim et al., 2024a) using LLaMA-3.2-3B-Instruct as the base model (see Appendix D), and deploy it using vLLM for optimized runtime performance. We present our hyperparameter study of each component in Appendix J. Knowledge Sources. We employ faiss (Douze et al., 2024) for efficient vector searching over the dense index. Following the Self-RAG (Asai et al., 2023), we utilize the Wikipedia 2018 (Izacard et al., 2023) by default, while specifically using the Wikipedia 2020 for PopQA to access more recent information. To optimize retrieval efficiency, we partition the index into five segments.